OpenStudy (anonymous):

Transform each polar equation to an equation in rectangular coordinates and identify its shape. a. θ = -(π / 6) b. r = (4 / (2cosθ - 3sinθ))

OpenStudy (anonymous):

@dan815 @freckles @Luigi0210 @radar @rvc @sleepyjess

OpenStudy (anonymous):

I think for A, r = 2 but I'm not sure

OpenStudy (freckles):

\[\theta=\frac{-\pi}{6}\] actually describes a line